OpenStudy (anonymous):

A worker can build a fence in 8 hours, With the help of an assistant, the fence can be built in 5 hours. How long should it take the assistant, working alone, to build the fence?

OpenStudy (anonymous):

Wouldn't it be the same amount of hours?

OpenStudy (anonymous):

What was wrong with my previous answer?

OpenStudy (anonymous):

9 hours.

OpenStudy (anonymous):

Not necessarily. And could you please show some work?

OpenStudy (anonymous):

well lets say the assistant worked at the same rate as the worker. then it would take half the time, 4 hours right.

OpenStudy (anonymous):

Oh. Nevermind. The answer is 40/3 hours.

OpenStudy (anonymous):

oh :(

OpenStudy (anonymous):

13 hours? how?

OpenStudy (anonymous):

Worker=W Assistant=A 1/A+1/W=1/5 1/A+1/8=1/5 8+A/8A=1/5 5*(8+A)=8A 40+5A=8A 3A=40 A=40/3 hours

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

why 1/A ?

OpenStudy (anonymous):

It's the work rate.

OpenStudy (anonymous):

Algebra.

OpenStudy (anonymous):

why is the rate 1/A ?

OpenStudy (anonymous):

sorry, i dont get it

OpenStudy (anonymous):

So it takes you 6 minutes to shower, right?

OpenStudy (anonymous):

You've showered for 2 minutes already. Doesn't that mean that you showered for 2/6 of the regular shower time or 1/3?

OpenStudy (anonymous):

http://www.purplemath.com/modules/workprob.htm

OpenStudy (anonymous):

That makes sense.